# Guardian: Detecting Robotic Planning and Execution Errors with Vision-Language Models

### v2

@@ Abstract
Robust robotic manipulation requires reliable failure detection and recovery. Although current Vision-Language Models (VLMs) show promise, their accuracy and generalization are limited by the scarcity of failure data. To address this data gap, we propose an automatic robot failure synthesis approach that procedurally perturbs successful trajectories to generate diverse planning and execution failures. This method produces not only binary classification labels but also fine-grained failure categories and step-by-step reasoning traces in both simulation and the real world. With it, we construct three new failure detection benchmarks: RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail, substantially expanding the diversity and scale of existing failure datasets. We then train Guardian, a VLM with multi-view images for detailed failure reasoning and detection. Guardian achieves state-of-the-art performance on both existing and newly introduced benchmarks. It also effectively improves task success rates when integrated into a state-of-the-art manipulation system in simulation and real robots, demonstrating the impact of our generated failure data. Code, Data, and Models available at https://www.di.ens.fr/willow/research/guardian/.
@@ I Introduction
Despite recent advances in task planning and action execution enabled by Large Language Models (LLMs) [mistralsmall, grattafiori2024llama3herdmodels] and Vision-Language Models (VLMs) [zhu2025internvl3exploringadvancedtraining], current robotic manipulation systems remain vulnerable to a wide range of failures [sinha2023systemlevelviewoutofdistributiondata, Kawaharazuka_2024, kroemer2020reviewrobotlearningmanipulation]. They can suffer from incorrect task plans [huang2022language] due to LLM hallucinations [Huang_2025], perception errors such as confusing similar objects  [garcia2025generalizablevisionlanguageroboticmanipulation, goyal2023rvtroboticviewtransformer], and low-level control issues such as unstable grasps or slippage [wu2025robomind]. These failures often compound during task execution, severely hindering the reliability of robots in real-world scenarios. To improve robustness, there is a growing interest in models that can leverage vision and language to reason about task execution and detect robot failures [liu2023reflect, chen2024automatingrobotfailurerecovery, duan2025aha].
A key challenge in robotic failure detection is the lack of comprehensive failure datasets. Most real-world robot datasets consist of successful demonstrations [khazatsky2025droidlargescaleinthewildrobot, embodimentcollaboration2024openxembodimentroboticlearning, pumacay2024colosseumbenchmarkevaluatinggeneralization], offering limited insight into failure modes. Automatically collecting failures by running trained policies is slow and unsafe, while manually collecting failure data [chen2024automatingrobotfailurerecovery, bu2025agibot] is labor-intensive and often lacks diversity and realism. As a result, prior work [duan2025aha] largely relies on simulation to construct failure cases, which however suffers from the notorious sim-to-real gap [simtorealgapzhao2020]. Furthermore, they predominantly focus on execution errors, with limited coverage of planning failures [duan2025aha] and lack fine-grained failure reasons [liu2023reflect, agia2024unpackingfailuremodesgenerative, duan2025aha, chen2024automatingrobotfailurerecovery, bu2025agibot]. This scarcity of diverse, realistic, and richly annotated failure data significantly hampers the development and evaluation of failure detection methods.
Another challenge is effectively leveraging and reasoning over visual context for failure detection. REFLECT [liu2023reflect] converts images into textual descriptions and then leverages LLM reasoning to detect failures, but this multi-stage pipeline is vulnerable to error accumulation from incomplete or inaccurate image captions. Other recent approaches [agia2024unpackingfailuremodesgenerative, etukuru2024robotutilitymodelsgeneral] use VLMs to process images directly, yet typically rely on single-view observations, making them susceptible to occlusions and incomplete scene understanding. AHA [duan2025aha] addresses this by concatenating multi-view images across timesteps into a single grid-based representation; although this expands visual coverage, the compressed format can hinder fine-grained spatial-temporal understanding. Moreover, most existing methods output categorical labels without providing explicit reasoning traces, limiting interpretability and the ability to diagnose complex failure modes.
Figure 1: Illustration of our Guardian model - a VLM fine-tuned on our constructed failure datasets. It detects planning failures (top) and execution failures (bottom) in robotic manipulation.
To address these challenges, we introduce an automatic failure generation approach that produces diverse planning and execution failures, and we train a reasoning VLM Guardian on the resulting data for robust failure detection. Our method procedurally perturbs successful robot trajectories in both simulation and real-world settings to generate failures annotated with fine-grained categories and step-by-step reasoning traces. Using this pipeline, we construct three new datasets: RLBench-Fail (14K execution and 7K planning samples in the RLBench simulator [james2019rlbenchrobotlearningbenchmark]), BridgeDataV2-Fail (10K execution and 6K planning samples using the real-robot BridgeDataV2 data [walke2024bridgedatav2datasetrobot]), and UR5-Fail (570 execution and 370 planning samples), our collected real-robot failure dataset with policy-driven rollouts. We then develop Guardian, a reasoning VLM fine-tuned for failure detection. As illustrated in Fig. 1, Guardian formulates failure detection as a visual question-answering task, reasoning over task instructions, subtask descriptions, and the visual scene. It leverages high-resolution, multi-view images to provide fine-grained and interpretable failure categorization. Guardian achieves state-of-the-art performance on both RLBench-Fail and BridgeDataV2-Fail, and generalizes in a zero-shot manner to RoboFail [liu2023reflect] and our UR5-Fail dataset. We further demonstrate Guardian’s plug-and-play capability as a feedback module by integrating it into a robotic manipulation system 3D-LOTUS++ [garcia2025generalizablevisionlanguageroboticmanipulation]. Guardian significantly improves task success through planning and execution monitoring in both simulated and real-world tasks.
@@
  • We propose a novel robot failure generation approach that perturbs successful robot demonstrations in both simulation and real-world settings, and autonomously annotates failures with fine-grained categories and reasoning traces. This yields three new datasets for robot failure detection: RLBench-Fail, BridgeDataV2-Fail, and UR5-Fail.
  • We introduce Guardian, a fine-tuned VLM for planning and execution failure detection. Guardian integrates high-resolution, multi-view visual inputs with explicit reasoning to predict fine-grained categorization.
  • Guardian achieves state-of-the-art performance across four benchmarks, covering both in- and out-of-domain datasets. It further improves task completion of a vision-language manipulation framework in simulated and real robot tasks.
@@ II Related Work
Robot failure datasets. Existing robot failure datasets cover only limited settings. RoboFail [liu2023reflect] provides a small collection of hand-crafted simulation and real-world failures. Sentinel [agia2024unpackingfailuremodesgenerative] generates failures by rolling out a trained policy on out-of-distribution scenarios created through randomizing object scales and poses, but it includes only four tasks. AHA [duan2025aha] perturbs trajectories in RLBench simulator [james2019rlbenchrobotlearningbenchmark] to create over 49K image–query pairs, yet the data is restricted to simulation and does not include high-level planning failures. Moreover, none of these datasets provides step-by-step reasoning annotations, limiting the development of models capable of fine-grained failure reasoning. In this work, we propose an automated approach that generates large-scale datasets with diverse planning and execution failures, enriched with fine-grained categories and reasoning traces in both simulation and real-robot environments.
Failure detection in robotic manipulation. Traditional monitoring methods rely on explicit models of tasks, identifying model deviations [de1998execution, gianni2011unified]. Unlike LLM-based approaches, they depend on rigid model-based predictions. Recent works formulate failure detection as a Question Answering (QA) task, by zero-shot prompting LLMs and VLMs [agia2024unpackingfailuremodesgenerative, etukuru2024robotutilitymodelsgeneral, chen2024automatingrobotfailurerecovery, ma2024generative, shirai2024visionlanguageinterpreterrobottask, skreta2024replanroboticreplanningperception, mei2024replanvlmreplanningrobotictasks]. For instance, REFLECT [liu2023reflect] summarizes multi-sensory inputs as texts, including audio, and feeds the text to an LLM, but its reliance on consecutive off-the-shelf models leads to accumulated errors. Consequently, efforts have been made to fine-tune VLMs specifically for robotic failure detection. SuccessVQA [pmlr-v232-du23b] fine-tunes Flamingo [flamingo] using both simulated and real-world data, but evaluates only the final task success, while Guardian is designed for more granular monitoring, checking the execution of each subtask. Closest to our work, AHA [duan2025aha] fine-tunes LLaVA-1.5 [liu2023visualinstructiontuning] on multi-class failures generated in RLBench. AHA concatenates multiple camera viewpoints and keysteps into a single image, leading to compressed visual inputs. Unlike AHA and SuccessVQA, which treat failure detection as direct classification, Sentinel [agia2024unpackingfailuremodesgenerative] and Cosmos-Reason1 [nvidia2025cosmosreason1physicalcommonsense] frame it as a reason-first process with explicit chain-of-thought (CoT) [10.5555/3600270.3602070] reasoning but focus only on single-view binary failure detection. Sentinel proposes a zero-shot CoT prompting framework to leverage off-the-shelf VLMs, while Cosmos-Reason1 is fine-tuned for physical reasoning. Extending prior approaches, Guardian unifies planning and execution verification via chain-of-thought reasoning over separated multi-view inputs, for multi-class failure detection.
Learning-based robotic manipulation policies. Recent robotic policies have achieved remarkable performance on manipulation tasks [garcia2025generalizablevisionlanguageroboticmanipulation, embodimentcollaboration2024openxembodimentroboticlearning, ze20243ddiffusionpolicygeneralizable, black2025pi]. However, their real-world applicability is still limited by planning and execution errors leading to task failures [wu2025robomind]. Inspired by Manipulate-Anything [duan2024manipulateanythingautomatingrealworldrobots] and Robot Utility Models [etukuru2024robotutilitymodelsgeneral], which demonstrated significant benefits from integrating VLM verifiers into their pipelines, Guardian is designed as a plug-and-play verification module trained on specific failure data. It can seamlessly integrate into existing policies, enhancing robustness by verifying planning and execution stages.
III Robot Failure Datasets Construction
Figure 2: Failure Data Generation Pipeline. We introduce a novel generation pipeline generating failure cases both online in simulation (RLBench), and offline on the real-world dataset (BridgeDataV2). For each positive example, given its correct plan and successful trajectory, we generate a corresponding incorrect plan and unsuccessful trajectory.
@@ III-A Data Sources
Simulated data enables controlled failure generation through procedural perturbations [duan2025aha], while real robot data reduces the sim-to-real gap but requires substantial human supervision [liu2023reflect]. To balance precise control and real-world fidelity, we use both simulated and real-robot datasets to construct robot failure datasets. We propose an automated method that derives planning and execution failures directly from successful demonstrations, avoiding manual failure collection. In both domains, tasks are decomposed into subtasks with corresponding video segments, which form the basis for generating failures. Fig. 2 (middle) illustrates successful episodes from the simulated and real robot datasets.
Simulated Data. We use the RLBench [james2019rlbenchrobotlearningbenchmark] simulator, selecting 52 tasks from RLBench-18Task [shridhar2022perceiveractormultitasktransformerrobotic] and GemBench [garcia2025generalizablevisionlanguageroboticmanipulation] benchmarks in our training data. For each task, we generate successful scripted trajectories with varied object placements and segment them into subtasks following 3D-LOTUS++ [garcia2025generalizablevisionlanguageroboticmanipulation].
@@ III-B Automated Failure Data Generation
Planning Failures. As shown in Fig. 2 (top), we construct five types of planning failures:
@@
    Imprecise grasping/pushing – the gripper attempts to grasp or push the correct object by moving toward it and closing its jaws, but misses due to inaccurate positioning.
Figure 3: Training examples. Guardian is trained on reasoning traces produced by our automatic data generation pipeline.
@@ III-C Chain-of-Thought (CoT) Generation
TABLE I: Robot failure datasets sizes across splits.
Training Validation Test Dataset Exec Plan Exec Plan Exec Plan RoboFail [liu2023reflect] - - - - 153 30 RLBench-Fail 12358 5808 1000 500 1000 500 BridgeDataV2-Fail 7830 4880 1000 500 1000 500 UR5-Fail 400 200 30 30 140 140
@@ III-D Real-Robot, Policy-Driven Data Collection
We curate UR5-Fail, a real-robot dataset, collected using a UR5 arm with three cameras. We run the 3D-LOTUS++ policy [garcia2025generalizablevisionlanguageroboticmanipulation] on 34 tasks, recording initial and final multi-view images for each subtask. Subtasks are manually labeled as success or failure to obtain execution failure data. For planning failures, we annotate ground-truth plans and generate failures using the method described in Sec. III-B. Unlike RoboFail [liu2023reflect], which is single-view and relies solely on teleoperation, UR5-Fail is three-view and features autonomous policy rollouts yielding more realistic failures.

### v3

# Scaling Cross-Environment Failure Reasoning Data for Vision-Language Robotic Manipulation

@@ Abstract
Robust robotic manipulation requires reliable failure detection and recovery. Although recent Vision-Language Models (VLMs) show promise in robot failure detection, their generalization is severely limited by the scarcity and narrow coverage of failure data. To address this bottleneck, we propose an automatic framework for generating diverse robotic planning and execution failures across both simulated and real-world environments. Our approach perturbs successful manipulation trajectories to synthesize failures that reflect realistic failure distributions, and leverages VLMs to produce structured step-by-step reasoning traces. This yields FailCoT, a large-scale failure reasoning dataset built upon the RLBench simulator and the BridgeDataV2 real-robot dataset. Using FailCoT, we train Guardian, a multi-view reasoning VLM for unified planning and execution verification. Guardian achieves state-of-the-art performance on three unseen real-world benchmarks: RoboFail, RoboVQA, and our newly introduced UR5-Fail. When integrated with a state-of-the-art LLM-based manipulation policy, it consistently boosts task success rates in both simulation and real-world deployment. These results demonstrate that scaling high-quality failure reasoning data is critical for improving generalization in robotic failure detection. Code, Data, and Models available at https://www.di.ens.fr/willow/research/guardian/.
@@ I Introduction
Recent advances in Large Language Models (LLMs) [mistralsmall, grattafiori2024llama3herdmodels] and Vision-Language Models (VLMs) [zhu2025internvl3exploringadvancedtraining] have significantly improved vision-language robotic manipulation. Nevertheless, existing models remain vulnerable to diverse failures [sinha2023systemlevelviewoutofdistributiondata, Kawaharazuka_2024, kroemer2020reviewrobotlearningmanipulation] such as incorrect task decomposition, object confusion, or unstable grasps, which compound over long horizons and degrade real-world reliability. As a result, automatic failure detection and recovery has received growing research attention [liu2023reflect, chen2024automatingrobotfailurerecovery, duan2025aha, agia2024unpackingfailuremodesgenerative, etukuru2024robotutilitymodelsgeneral, ifailsense2026, zeng2025vifailback].
Leveraging their strong generalization ability, LLMs and VLMs have been increasingly explored for failure detection. Some methods [liu2023reflect, etukuru2024robotutilitymodelsgeneral, duan2024manipulateanythingautomatingrealworldrobots], directly prompt pretrained foundation models to detect failures, optionally enhanced with chain-of-thought (CoT) reasoning [agia2024unpackingfailuremodesgenerative, nvidia2026cosmosreason2, cot2022] or multi-agent code generation [zhou2024code]. While promising, these approaches suffer from a large domain gap: robotic observations differ substantially from web-scale pretraining data, and accurate failure detection requires fine-grained, embodied reasoning beyond generic visual understanding. Therefore, recent work [duan2025aha, ifailsense2026, robofac2025] has shifted toward fine-tuning VLMs on robot failure datasets to better bridge the gap.
A fundamental bottleneck, however, is the scarcity of large-scale, high-quality failure data. Most robot learning datasets predominantly contain successful demonstrations [khazatsky2025droidlargescaleinthewildrobot, embodimentcollaboration2024openxembodimentroboticlearning, pumacay2024colosseumbenchmarkevaluatinggeneralization], providing limited failure examples. Collecting failures by rolling out policies is time-consuming and potentially unsafe, while manual curation [chen2024automatingrobotfailurerecovery, bu2025agibot] is labor-intensive and typically lacks diversity. Several prior approaches [duan2025aha, agia2024unpackingfailuremodesgenerative, dai2024racer] rely on simulated failure examples, but these suffer from sim-to-real gap [simtorealgapzhao2020], and provide limited coverage of both low-level execution errors and high-level planning failures [ifailsense2026].
To address these limitations, we propose an automatic failure generation framework that synthesizes diverse planning and execution failures across simulated and real-world environments. Starting from successful demonstrations, we procedurally perturb task plans and subtask executions to create realistic failures, augmenting each example with structured step-by-step reasoning traces. This enables the construction of FailCoT, a large-scale failure reasoning dataset containing over 30K training examples. It includes RLBench-Fail built using the RLBench simulator [james2019rlbenchrobotlearningbenchmark] and BridgeDataV2-Fail derived from the BridgeDataV2 real-robot dataset [walke2024bridgedatav2datasetrobot], see Figure 1. FailCoT provides balanced success and failure samples, multi-view visual observations, and explicit CoT supervision for plan and subtask-level verification.
Building on FailCoT, we develop Guardian, a multi-view reasoning VLM fine-tuned for unified planning and execution failure detection. Guardian formulates verification as a visual question answering problem: conditioned on task instructions, proposed plans or subtasks, and multi-view observations, it produces explicit reasoning traces to enhance predicting failures. To further support realistic evaluation, we introduce a new real-robot benchmark UR5-Fail constructed using the same failure generation framework. Guardian achieves state-of-the-art performance on three unseen real-world failure benchmarks, namely RoboFail [liu2023reflect], RoboVQA [sermanet2024] and UR5-Fail. When integrated as a plug-and-play verification module into a LLM-based manipulation system, Guardian improves task success in both simulation and real-robot experiments. Extensive ablations demonstrate the benefits of scaling structured, cross-environment failure reasoning data.
@@
  • We propose an automatic cross-environment failure synthesis framework that generates diverse planning and execution errors with structured reasoning supervision, resulting in the large-scale robot failure dataset FailCoT.
  • We develop Guardian, a multi-view reasoning VLM fine-tuned on the FailCoT dataset for unified planning and execution failure detection.
  • We show that scaling structured failure reasoning data yields state-of-the-art detection performance and improves task success when deployed as a plug-and-play verifier.
@@
Figure 1: Failure Data Generation Framework. We construct failure cases both online in simulation (RLBench), and offline on the real-world dataset (BridgeDataV2). For each positive example, given its correct plan and successful trajectory, we generate a corresponding incorrect plan and unsuccessful trajectory.
@@ II Related Work
Vision-Language Robotic Manipulation. Recent advances in foundation models [zhu2025internvl3exploringadvancedtraining, openai2024gpt4ocard] have significantly improved vision–language robotic manipulation. End-to-end vision–language–action (VLA) policies such as Gr00T [groot2025] and $\pi_{0}$ [black2024pi] directly predict action sequences from 2D images and task instructions. To enhance spatial reasoning, 3D-based VLAs have further been proposed [garcia2025generalizablevisionlanguageroboticmanipulation, goyal2023rvtroboticviewtransformer]. Notably, 3D-LOTUS++ [garcia2025generalizablevisionlanguageroboticmanipulation] achieves state-of-the-art performance on challenging generalizable manipulation tasks through a modular design that combines LLM-based task planning, visual grounding modules, and 3D-based execution policies. Despite this progress, robust robotic manipulation remains challenging, as planning mistakes and execution errors accumulate over long horizons [wu2025robomind]. To improve robustness, recent approaches incorporate runtime monitoring of failures to trigger policy correction or retry [duan2024manipulateanythingautomatingrealworldrobots, etukuru2024robotutilitymodelsgeneral, dai2024racer]. In this work, we advance automatic failure detection by scaling structured training data and demonstrate improved integration with manipulation policies.
Robotic Failure Detection Methods. Early rule-based failure detection methods [de1998execution, gianni2011unified] struggle to generalize beyond predefined task structures. Learning-based approaches address this limitation and can be broadly categorized based on whether they require robot failure data for training.
Training-free methods. One line of work performs out-of-distribution (OOD) detection [canwedetectfailurewithoutfailuredata2025] or temporal inconsistency detection [agia2024unpackingfailuremodesgenerative] over internal policy representations to flag failures without failure supervision. Another line prompts LLMs or VLMs for failure assessment [ahmad2025unifiedframework] using techniques such as hierarchical CoT reasoning [liu2023reflect, agia2024unpackingfailuremodesgenerative] or constraint-aware visual programming [zhou2024code]. While these approaches avoid collecting robot failure data, they rely on policy-specific signals or prompt engineering, and cannot learn from robotic failure data for better accuracy.
Training-based methods. Recent works [duan2025aha, ifailsense2026, nvidia2026cosmosreason2, gu2025safe, pmlr-v232-du23b, armor2026] fine-tune VLMs as failure detectors using annotated trajectories, but mostly target execution-stage verification. SuccessVQA [pmlr-v232-du23b] performs coarse task-level success assessment. AHA [duan2025aha] and I-Fail-Sense [ifailsense2026] compress multi-view inputs into a single concatenated image for detecting subtask-level execution failures or instruction–behavior misalignment. SAFE [gu2025safe] relies on internal policy representation to train a failure classifier. Cosmos-Reason [nvidia2026cosmosreason2] addresses general embodied reasoning trained with supervised fine-tuning and reinforcement learning, where failure detection is one downstream task. ARMOR [armor2026] uses multi-round self-refinement with separate detection and reasoning heads, but it trains only on post-execution failures and requires multiple inference passes per sample. Compared to prior methods, our Guardian model leverages large-scale failure reasoning data to enable multi-view, explicit reasoning for unified planning and execution verification, achieving state-of-the-art performance.
Robot Failure Datasets. Collecting real-world robot failures at scale is challenging: policy rollouts are time-consuming, potentially unsafe, and require extensive manual annotation. RoboFail [liu2023reflect] provides a hand-crafted dataset spanning simulation and real settings, but covers limited tasks and failure modes. ViFailback [zeng2025vifailback] focuses on single-view, single-embodiment real-world diagnosis and requires substantial teleoperation effort. To reduce collection cost, several works rely on synthetic failure generation. Sentinel [agia2024unpackingfailuremodesgenerative] induces failures via out-of-distribution rollouts but covers only four tasks. SAFE [gu2025safe] utilizes the final sparse reward for policy rollouts and lacks dense failure supervision for each step. AHA [duan2025aha] perturbs trajectories in RLBench [james2019rlbenchrobotlearningbenchmark], generating large-scale purely simulated data, yet excludes high-level planning failures, and the dataset has not been publicly released. RoboFAC [robofac2025] adds reasoning annotations in simulation, with only a limited real-world subset via manual teleoperation. I-Fail-Sense [ifailsense2026] synthesizes failures from RLBench and DROID, but focuses primarily on semantic mismatches, leaving planning and low-level control errors underexplored. In contrast, we propose an automated pipeline that generates diverse planning and execution failures across both simulation and real robots, producing realistic failure modes at scale with multi-view observations and fine-grained, step-by-step reasoning supervision.
III FailCoT: Cross-Environment Robot Failure Reasoning Datasets
@@ III-A Data Sources
Simulated data enables controlled failure generation through procedural perturbations [duan2025aha], while real robot data reduces the sim-to-real gap but requires substantial human supervision [liu2023reflect]. To balance precise control and real-world fidelity, we use both simulated and real-robot datasets to construct robot failure datasets. We propose an automated method that derives planning and execution failures directly from successful demonstrations, avoiding manual failure collection. In both domains, tasks are decomposed into subtasks with corresponding video segments, which form the basis for generating failures. Fig. 1 (middle row) illustrates successful episodes from the simulated and real robot datasets.
Simulated Data. We use the RLBench [james2019rlbenchrobotlearningbenchmark] simulator, selecting 52 tasks from RLBench-18Task [shridhar2022perceiveractormultitasktransformerrobotic] and GemBench [garcia2025generalizablevisionlanguageroboticmanipulation] benchmarks in our training data. For each task, we generate successful scripted trajectories with varied object placements and segment them into subtasks following [garcia2025generalizablevisionlanguageroboticmanipulation].
@@ III-B Automated Failure Data Generation
Planning Failures. As shown in Fig. 1 (top row), we construct five types of planning failures:
@@
    Imprecise grasping/pushing – the gripper attempts to grasp or push the correct object by moving toward it and closing its jaws, but misses it due to inaccurate positioning.
@@ III-D Real-Robot, Policy-Driven Data Collection
To further support realistic evaluation, we curate UR5-Fail, a real-robot dataset, collected using a UR5 arm with three cameras. We run the 3D-LOTUS++ policy [garcia2025generalizablevisionlanguageroboticmanipulation] on 16 unique tasks, recording initial and final multi-view images for each subtask. Subtasks are manually labeled as success or failure to obtain execution failure data. For planning failures, we annotate ground-truth plans and generate failures using the method described in Sec. III-B. Unlike RoboFail [liu2023reflect], which is single-view and relies solely on teleoperation, UR5-Fail is three-view and features autonomous policy rollouts yielding more realistic failures.
III-E Dataset Statistics and Evaluation
FailCoT (RLBench-Fail, BridgeDataV2-Fail) and UR5-Fail, contain balanced success/failure examples across both planning and execution, with reasoning traces. FailCoT is split into training, validation, and test sets, with the validation and test sets featuring unseen tasks/environments to evaluate generalization, see Table I top. Table I bottom compares UR5-Fail with two existing real-world datasets and shows a more balanced distribution between execution and planning.
TABLE I: Statistics of the FailCoT dataset and three real-world robot failure detection benchmarks. Our constructed datasets (FailCoT and UR5-Fail) contain balanced success and failure cases covering both execution and planning errors.
Training Validation Test
Dataset Env. Exec Plan Exec Plan Exec Plan
FailCoT (Ours)
   RLBench-Fail Sim 12358 5808 1000 500 1000 500
   BridgeDataV2-Fail Real 7830 4880 1000 500 1000 500
Real-World Robot Failure Detection Benchmarks
UR5-Fail (Ours) Real - - - - 140 140
RoboFail [liu2023reflect] Real - - - - 153 30
RoboVQA [sermanet2024] Real - - - - 357 -